# Tensor Decomposition Networks for Fast Machine Learning Interatomic Potential Computations

### v1

@@ Abstract
$\rm{SO}(3)$-equivariant networks are the dominant models for machine learning interatomic potentials (MLIPs). The key operation of such networks is the Clebsch-Gordan (CG) tensor product, which is computationally expensive. To accelerate the computation, we develop tensor decomposition networks (TDNs) as a class of approximately equivariant networks whose CG tensor products are replaced by low-rank tensor decompositions, such as the CANDECOMP/PARAFAC (CP) decomposition. With the CP decomposition, we prove (i) a uniform bound on the induced error of $\rm{SO}(3)$-equivariance, and (ii) the universality of approximating any equivariant bilinear map. To further reduce the number of parameters, we propose path-weight sharing that ties all multiplicity-space weights across the $\mathcal{O}(L^{3})$ CG paths into a single path without compromising equivariance, where $L$ is the maximum angular degree. The resulting layer acts as a plug-and-play replacement for tensor products in existing networks, and the computational complexity of tensor products is reduced from $\mathcal{O}(L^{6})$ to $\mathcal{O}(L^{4})$. We evaluate TDNs on PubChemQCR, a newly curated molecular relaxation dataset containing 105 million DFT-calculated snapshots. We also use existing datasets, including OC20, and OC22. Results show that TDNs achieve competitive performance with dramatic speedup in computations.
@@ 2 Introduction
Symmetry is a fundamental aspect of molecular and material systems (zhang2023artificial, ), making it a crucial consideration in developing machine learning interatomic potentials (MLIPs). Equivariant graph neural networks have emerged as dominant frameworks in this domain. Usually, equivariant models incorporate directional features and spherical harmonics to maintain equivariance under rotation symmetry (gasteiger2019directional, ; satorras2021n, ; schutt2021equivariant, ). Among these, tensor product (TP) operations play a central role in fusing equivariant features, providing a powerful mechanism for building expressive models that adhere to $\rm{SO}(3)$-equivariance (thomas2018tensor, ; liao2022equiformer, ; batatia2022mace, ).
However, the computational cost of tensor product operations grows rapidly with the maximum angular degree $L$, reaching $\mathcal{O}(L^{6})$ in conventional implementations. Recent efforts to mitigate this cost have focused on accelerating the TP operation or applying frame averaging (FA) to enforce equivariance (duval2023faenet, ; dym2024equivariant, ; lin2024equivariance, ). While frame averaging is architecture-agnostic, it suffers from discontinuity issues (dym2024equivariant, ). On the other hand, TP acceleration techniques, such as SO(2)-based convolutions and fast spherical Fourier transformations, reduce complexity but still involve substantial parameter costs and computational overhead. Thus, there remains a gap in developing a method that simultaneously reduces computational complexity and parameter count while maintaining high accuracy and equivariance.
In this work, we propose Tensor Decomposition Networks (TDNs), a new class of approximately equivariant networks that replace the standard CG tensor product with low-rank tensor decompositions based on CANDECOMP/PARAFAC (CP) decomposition. TDNs introduce a CP decomposition that provides error bounds on equivariance and universality, ensuring consistency under $\rm{SO}(3)$ transformations. Additionally, a path-weight sharing mechanism consolidates multiplicity-space weights across CG paths, significantly reducing the parameter count from $\mathcal{O}(c^{2}L^{3})$ to $\mathcal{O}(c^{2})$ with $c$ the multiplicity. The resulting layer maintains the expressive power of conventional TP while lowering computational complexity from $\mathcal{O}(L^{6})$ to $\mathcal{O}(L^{4})$. We validate TDNs on a newly curated relaxation dataset with 105 million DFT-calculated molecular snapshots, along with the established OC20 and OC22 datasets, demonstrating competitive accuracy with substantial computational speedup.
@@ 3 Preliminaries and Related Work
Tensor Product Acceleration. Among these equivariant networks, tensor product (thomas2018tensor, ; fuchs2020se, ; liao2022equiformer, ; batzner20223, ; batatia2022mace, ; unke2021se, ; yu2023efficient, ) is one of the most important components that fuse two equivariant features into one. It provides a powerful and expressive way (dym2020universality, ) to build equivariant networks, while the computational cost of TP is usually considerable. Therefore, there are several directions to accelerate the equivariant networks. First direction is to accelerate TP. eSCN (liao2024equiformerv, ; passaro2023reducing, ) proposes to reduce the SO(3) convolution into SO(2) for TP when one of the inputs of TP is spherical harmonics. Gaunt tensor product (luoenabling, ) makes use of fast spherical Fourier transformation to perform the TP. The other direction is to apply frame averaging (FA) (duval2023faenet, ; dym2024equivariant, ; lin2024equivariance, ), which uses group elements from an equivariant set-value function called frame to transform the input data and subsequently the model’s output, enabling any models to obtain the desired symmetries. Although it is flexible and has no requirement for model architectures, it faces an unsolvable discontinuity problem (dym2024equivariant, ).
@@ 4 Tensor Decomposition Networks
This section presents the techniques employed in the proposed Tensor Decomposition Network (TDN). In Section 4.1, we introduce a CP-decomposition-based approximation for the tensor product, and in Section 4.2, we detail a path-wise weight-sharing scheme. Together, these strategies effectively reduce both computational cost and parameter count. In Section 4.3, we analyze the computational complexity of the approximate tensor product, and in Section 4.4, we discuss its error bound and universality.
@@ 4.1 Tensor Product and Its Approximation
Before developing our tensor-product approximation we recall the canonical definition of the tensor product in the simplest non-trivial two-order case. The multi-order counterpart and its CP decomposition are introduced in Appendix A. In practice, higher-rank tensors are stored as flattened vectors via a fixed index ordering; this reshaping preserves the vector-space operations. Without loss of generality, we present the following tensor product definition. Let $V_{1}=\mathbb{R}^{d_{1}}$, $V_{2}=\mathbb{R}^{d_{2}}$, and $V_{3}=\mathbb{R}^{d_{3}}$ be finite-dimensional real vector spaces with fixed ordered bases $\{{\bm{e}}_{i}\}_{i=1}^{d_{1}},\{{\bm{f}}_{j}\}_{j=1}^{d_{2}},\{{\bm{g}}_{k}\}_{k=1}^{d_{3}}$. The tensor product $V_{1}\otimes V_{2}$ is the space that corepresents bilinear maps: for every bilinear $m\colon V_{1}\times V_{2}\to V_{3}$ there exists a unique linear map $\widetilde{m}\colon V_{1}\otimes V_{2}\;\longrightarrow\;V_{3}$ such that $m({\bm{x}},{\bm{y}})=\widetilde{m}({\bm{x}}\otimes{\bm{y}})$. With respect to the chosen bases, $\widetilde{m}$ is encoded by a three–way tensor
@@
where ${\bm{A}}\in\mathbb{R}^{d_{3}\times R}$, ${\bm{B}}\in\mathbb{R}^{d_{1}\times R}$, and ${\bm{C}}\in\mathbb{R}^{d_{2}\times R}$ are factor matrices capturing the modes of the tensor. Substituting Eq. 2 into Eq. 1 gives
@@
The key idea of CP decomposition is to break down the CG coefficient tensor ${\bm{M}}$ into low-rank matrices. We demonstrate a simple example of CP decomposition for TP with $\ell_{1},\ell_{2}=1$ and $\ell_{3}=(0,1,2)$, as shown in Fig. 1. Eq. 3 allows for the simple batch-wise use of the Hadamard product with a global hyperparameter $R$ instead of computing tensor product per path. The higher $R$, the more accurate the tensor product result is. For acceleration, we do not need full rank $R$ and we can use a small $R$ while keeping a low approximation error. We show the rank-error curve in Fig. 2.
@@ 4.2 Path-Weight Sharing Tensor Product
We first unify the multiplicities across all degrees $\ell$ by setting a common multiplicity $c$. Letting ${\bm{x}}^{(\ell)}\in\mathbb{R}^{c\times(2\ell+1)}$ denote the multiplicity-$c$ irrep of degree $\ell$, we concatenate the multiplicity axes into a single channel:
$\tilde{\bm{x}}=\operatorname{concat}_{\ell=0}^{L}{\bm{x}}^{(\ell)}\in\mathbb{R}^{c\times(L+1)^{2}}.$
All irreps now share a common multiplicity index, so the linear projection and CG contraction over irreps features are executed as batched matrix operations whose operands reside contiguously in memory; this yields coalesced memory access on GPUs and improved cache locality on CPUs.
@@
We further reduce the parameter count by applying a path-weight sharing scheme. Let ${\bm{W}}_{\ell_{1},\ell_{2}}^{\ell}\in\mathbb{R}^{c\times c}$ denote the multiplicity-space weight matrix associated with the path $(\ell_{1},\ell_{2},\ell_{3})$, where $0\leq\ell_{1},\ell_{2},\ell_{3}\leq L$ and $c$ is the channel dimension of each irrep block. Instead of assigning a unique matrix for each path, we set all such matrices equal to a single learnable parameter ${\bm{W}}\in\mathbb{R}^{c\times c}$, i.e. ${\bm{W}}_{\ell_{1},\ell_{2}}^{\ell}\equiv{\bm{W}}$ for every admissible path $(\ell_{1},\ell_{2},\ell_{3})$. This collapses the $\mathcal{O}(L^{3})$ distinct weight tensors of the naïve implementation into one, reducing the parameter count from $\mathcal{O}\bigl{(}c^{2}L^{3}\bigr{)}$ to $\mathcal{O}\bigl{(}c^{2}\bigr{)}$. Because the sharing operates exclusively on multiplicity indices, the irrep content of each block and hence full $\mathrm{SO}(3)$ equivariance is retained. The resulting layer therefore generalizes the classic CG tensor product while offering an orders-of-magnitude reduction in parameters.
@@ 4.3 Complexity Analysis of Approximate Tensor Product
In this section, we analyze the computational complexity of the proposed approximate tensor product using CP decomposition. We first investigate the rank of the tensor product. Let $\operatorname{rank}_{\mathrm{CP}}({\bm{M}})$ denote the CP rank of the three-way tensor ${\bm{M}}\in\mathbb{R}^{d_{3}\times d_{1}\times d_{2}}$, representing the minimal rank $R$ such that there is an equality for Eq. 2, i.e.
@@
By using CP decomposition, the computational cost for evaluating the approximate tensor product in Eq. 3 reduces to $\mathcal{O}\bigl{(}R(d_{1}+d_{2}+d_{3})\bigr{)}$, a significant reduction compared to the full tensor contraction cost of $\mathcal{O}(d_{1}d_{2}d_{3})$ using Eq. 1. Similarly, the parameter budget decreases from $\mathcal{O}(d_{1}d_{2}d_{3})$ to $\mathcal{O}(R(d_{1}+d_{2}+d_{3}))$, which provides substantial savings when $R$ is small.
For the approximation of the CG tensor product, where $d_{1},d_{2},d_{3}\propto\mathcal{O}(L^{2})$, the computational complexity further reduces to $\mathcal{O}(RL^{2})$. In our experiments, we select $R=7L^{2}$. The error curve for the CP decomposition with varying $R$ is discussed in Section 5.1.
@@ 4.4 Error Bound and Universality Analysis
The error bound and universality analysis provide theoretical guarantees for the CP decomposition of the tensor product in the proposed Tensor Decomposition Network (TDN). This section establishes the error bound for both the approximation and equivariance, demonstrating how the approximation error depends on the spectral tail of the tensor’s singular values. Additionally, we formalizes the universality property of the CP decomposition, showing that as the rank $R$ increases, the CP-decomposition-based tensor product can accurately approximate any $\mathrm{SO}(3)$-equivariant bilinear map, thereby preserving the expressive power of the tensor product while reducing computational complexity.
@@
Given a rank $R\leq\operatorname{rank}_{\mathrm{CP}}({\bm{M}})$, one seeks CP-decomposition-based approximation $\widehat{\bm{M}}$ by minimizing $\lVert{\bm{M}}-\widehat{\bm{M}}\rVert$ in a chosen norm, typically the Frobenius norm. Although the non-convex optimization may admit spurious local minima, the optimization of CP decomposition possesses an essentially unique best rank-$R$ approximation under Kruskal’s condition (kolda2009tensor, ), and modern alternating least-squares or gradient methods converge to it under mild coherence assumptions (yang2023global, ). For error estimation in our setting, we specialize to the CG tensor product, where we let all irreps have the same maximum degree and all dimensions equal such that $d=d_{1}=d_{2}=d_{3}$. Given the singular values $\sigma_{k}^{(n)}$ of the mode-$n$ matricization ${\bm{M}}_{(n)}$ of ${\bm{M}}$ and truncating each ${\bm{M}}_{(n)}$ to rank $R_{T}=\lceil R^{1/3}\rceil$, a priori approximation error bound (de2000multilinear, ) gives
@@
The following result bounds this error uniformly over $\rm{SO}(3)$ with the proof in Section B.1. {theorem}Equivariance Error Bound of CP Decompositionequivariance-error Let CG tensor ${\bm{M}}\in\mathbb{R}^{d\times d\times d}$ and $\widehat{{\bm{M}}}$ be the rank-$R$ CP-decomposition-based approximation obtained by Frobenius minimization. For any rotation ${\bm{R}}\in SO(3)$ and any bounded irreps ${\bm{x}},{\bm{y}}\in\mathbb{R}^{d},\lVert{\bm{x}}\rVert,\lVert{\bm{y}}\rVert\leq C$, we have
[…]
where $R_{T}=\lceil R^{1/3}\rceil$ and $\sigma_{k}^{(n)}$ is the $k$-th singular value of mode-$n$ matricization of ${\bm{M}}$. Empirical estimates of both the approximation and equivariance errors are provided in Section 5.1.
@@
Because the tensor product is universal for bilinear maps, any $\rm{SO}(3)$-equivariant bilinear operator can be expressed as a composition of a tensor product followed by a suitable ${\bm{M}}$ onto an equivariant subspace. Consequently, the CP-decomposition-based approximation developed above inherits this universality: for every equivariant tensor product there exists a rank-$R$ approximation that converges to it as $R\to\operatorname{rank}_{\mathrm{CP}}({\bm{M}})$. The following theorem formalizes the expressivity of our approximation scheme with the proof in Section B.2.
{theorem}
Universality of CP Decompositionuniversality Consider $\rm{SO}(3)$-irreps ${\bm{x}}\in V_{1}=\mathbb{R}^{d}$, ${\bm{y}}\in V_{2}=\mathbb{R}^{d_{2}}$ and $V_{3}=\mathbb{R}^{d_{3}}$. Fo any $\rm{SO}(3)$-equivariant bilinear map $\Phi$, there exist ${\bm{B}}\in\mathbb{R}^{d_{1}\times R}$, ${\bm{C}}\in\mathbb{R}^{d_{2}\times R}$, ${\bm{A}}\in\mathbb{R}^{R\times d_{3}}$ such that $\Phi$ can be written as
$\Phi({\bm{x}},{\bm{y}})={\bm{A}}({\bm{B}}^{\top}{\bm{x}}\odot{\bm{C}}^{\top}{\bm{y}})\in V_{3},$
if $R$ is sufficient large.

### v2

@@ Abstract
$\rm{SO}(3)$-equivariant networks are the dominant models for machine learning interatomic potentials (MLIPs). The key operation of such networks is the Clebsch-Gordan (CG) tensor product, which is computationally expensive. To accelerate the computation, we develop tensor decomposition networks (TDNs) as a class of approximately equivariant networks in which CG tensor products are replaced by low-rank tensor decompositions, such as the CANDECOMP/PARAFAC (CP) decomposition. With the CP decomposition, we prove (i) a uniform bound on the induced error of $\rm{SO}(3)$-equivariance, and (ii) the universality of approximating any equivariant bilinear map. To further reduce the number of parameters, we propose path-weight sharing that ties all multiplicity-space weights across the $\mathcal{O}(L^{3})$ CG paths into a single path without compromising equivariance, where $L$ is the maximum angular degree. The resulting layer acts as a plug-and-play replacement for tensor products in existing networks, and the computational complexity of tensor products is reduced from $\mathcal{O}(L^{6})$ to $\mathcal{O}(L^{4})$. We evaluate TDNs on PubChemQCR, a newly curated molecular relaxation dataset containing 105 million DFT-calculated snapshots. We also use existing datasets, including OC20, and OC22. Results show that TDNs achieve competitive performance with dramatic speedup in computations. Our code is publicly available as part of the AIRS library (https://github.com/divelab/AIRS/).
@@ 1 Introduction
Symmetry is a fundamental aspect of molecular and material systems (zhang2023artificial, ), making it a crucial consideration in developing machine learning interatomic potentials (MLIPs). Equivariant graph neural networks have emerged as dominant frameworks in this domain. Usually, equivariant models incorporate directional features and spherical harmonics to maintain equivariance under rotation symmetry (gasteiger2019directional, ; satorras2021n, ; schutt2021equivariant, ). Among these, tensor product (TP) operations play a central role in fusing equivariant features, providing a powerful mechanism for building expressive models that adhere to $\rm{SO}(3)$-equivariance (anderson2019cormorant, ; thomas2018tensor, ; liao2022equiformer, ; batatia2022mace, ).
However, the computational cost of tensor product operations grows rapidly with the maximum angular degree $L$, reaching $\mathcal{O}(L^{6})$ in conventional implementations. Recent efforts to mitigate this cost have focused on accelerating the TP operation or applying frame averaging (FA) to enforce equivariance (duval2023faenet, ; dym2024equivariant, ; lin2024equivariance, ). While frame averaging is architecture-agnostic, it suffers from discontinuity issues (dym2024equivariant, ). On the other hand, TP acceleration techniques, such as SO(2)-based convolutions (passaro2023reducing, ; liao2024equiformerv, ) and fast spherical Fourier transformations (luoenabling, ), reduce complexity but are no longer the standard CG tensor product with the same expressivity. Thus, there remains a gap in developing a method that simultaneously reduces computational complexity and parameter count while maintaining similar accuracy and equivariance of CG tensor product.
In this work, we propose Tensor Decomposition Networks (TDNs), a new class of approximately equivariant networks that replace the standard CG tensor product with low-rank tensor decompositions based on CANDECOMP/PARAFAC (CP) decomposition. TDNs introduce a CP decomposition that provides error bounds on equivariance and universality, ensuring consistency under $\rm{SO}(3)$ transformations. Additionally, a path-weight sharing mechanism consolidates multiplicity-space weights across CG paths, significantly reducing the parameter count from $\mathcal{O}(cL^{3})$ to $\mathcal{O}(c)$ with $c$ the parameter count of weight of each path. The resulting layer has a close expressive power of conventional TP while lowering computational complexity from $\mathcal{O}(L^{6})$ to $\mathcal{O}(L^{4})$. We validate TDNs on a newly curated relaxation dataset with 105 million DFT-calculated molecular snapshots, along with the established OC20 and OC22 datasets, demonstrating competitive accuracy with substantial computational speedup.
@@ 2 Preliminaries and Related Work
Tensor Product Acceleration. Among these equivariant networks, tensor product (anderson2019cormorant, ; thomas2018tensor, ; fuchs2020se, ; liao2022equiformer, ; batzner20223, ; batatia2022mace, ; unke2021se, ; yu2023efficient, ) is one of the most important components that fuse two equivariant features into one. It provides a powerful and expressive way (dym2020universality, ) to build equivariant networks, while the computational cost of TP is usually considerable. Therefore, there are several directions to accelerate the equivariant networks. First direction is to accelerate TP. eSCN (liao2024equiformerv, ; passaro2023reducing, ) proposes to reduce the SO(3) convolution into SO(2) for TP when one of the inputs of TP is spherical harmonics. Gaunt tensor product (luoenabling, ) makes use of fast spherical Fourier transformation to perform the TP. The other direction is to apply frame averaging (FA) (duval2023faenet, ; dym2024equivariant, ; lin2024equivariance, ), which uses group elements from an equivariant set-value function called frame to transform the input data and subsequently the model’s output, enabling any models to obtain the desired symmetries. Although it is flexible and has no requirement for model architectures, it faces an unsolvable discontinuity problem (dym2024equivariant, ).
@@ 3 Tensor Decomposition Networks
This section presents the techniques employed in the proposed Tensor Decomposition Network (TDN). In Section˜3.1, we introduce a CP-decomposition-based approximation for the tensor product, and in Section˜3.2, we detail a path-wise weight-sharing scheme. These strategies effectively reduce both computational cost and parameter count. In Section˜3.3, we analyze the computational complexity of the approximate tensor product, and in Section˜3.4, we discuss its error bound and universality.
@@ 3.1 Tensor Product and Its Approximation
Before developing our tensor‑product approximation we recall the canonical definition of the tensor product in the simplest non‑trivial two-order case. The multi‑order counterpart and its CP decomposition are introduced in Appendix˜A. In practice, higher‑rank tensors are stored as flattened vectors via a fixed index ordering; this reshaping preserves the vector‑space operations. Without loss of generality, we present the following tensor product definition. Let $V_{1}=\mathbb{R}^{d_{1}}$, $V_{2}=\mathbb{R}^{d_{2}}$, and $V_{3}=\mathbb{R}^{d_{3}}$ be finite-dimensional real vector spaces with ordered bases $\{{\bm{e}}_{i}\}_{i=1}^{d_{1}},\{{\bm{f}}_{j}\}_{j=1}^{d_{2}},\{{\bm{g}}_{k}\}_{k=1}^{d_{3}}$. The tensor product $V_{1}\otimes V_{2}$ is the space that corepresents bilinear maps: for every bilinear $m\colon V_{1}\times V_{2}\to V_{3}$ there exists a unique linear map $\widetilde{m}\colon V_{1}\otimes V_{2}\;\longrightarrow\;V_{3}$ such that $m({\bm{x}},{\bm{y}})=\widetilde{m}({\bm{x}}\otimes{\bm{y}})$. With respect to the chosen bases, $\widetilde{m}$ is encoded by a three–way tensor
@@
where ${\bm{A}}\in\mathbb{R}^{d_{3}\times R}$, ${\bm{B}}\in\mathbb{R}^{d_{1}\times R}$, and ${\bm{C}}\in\mathbb{R}^{d_{2}\times R}$ are factor matrices capturing the modes of the tensor. Substituting Eq.˜2 into Eq.˜1 gives
@@
The key idea of CP decomposition is to break down the CG coefficient tensor $\bm{M}$} into low-rank matrices. We demonstrate a simple example of CP decomposition for TP with $\ell_{1},\ell_{2}=1$ and $\ell_{3}\in\{0,1,2\}$, as shown in Fig.˜1. Eq.˜3 allows for the simple batch-wise use of the Hadamard product with a global hyperparameter rank $R$ instead of computing tensor product per path. The higher $R$, the more accurate the tensor product result is. For acceleration, we do not need full rank $R$ and we can use a small $R$ while keeping a low approximation error. We show the rank-error curve in Fig.˜2.
@@ 3.2 Path-Weight Sharing Tensor Product
We first unify the multiplicities across all degrees $\ell$ by setting a common multiplicity $d$. Letting ${\bm{x}}^{(\ell)}\in\mathbb{R}^{d\times(2\ell+1)}$ denote the multiplicity-$d$ irrep of degree $\ell$, we concatenate the multiplicity axes into a single channel:
$\tilde{\bm{x}}=\operatorname{concat}_{\ell=0}^{L}{\bm{x}}^{(\ell)}\in\mathbb{R}^{d\times(L+1)^{2}}.$
All irreps now share a common multiplicity index within, so the linear projection and CG contraction over irreps features are executed as batched matrix operations whose operands reside contiguously in memory; this yields coalesced memory access on GPUs and improved cache locality on CPUs.
@@
We further reduce the parameter count by applying a path-weight sharing scheme. Let ${\bm{W}}_{\ell_{1},\ell_{2}}^{\ell}$ denote the multiplicity-space weight matrix associated with the path $(\ell_{1},\ell_{2},\ell_{3})$, where $0\leq\ell_{1},\ell_{2},\ell_{3}\leq L$ and $c$ is the parameter count of weight of each path. Instead of assigning a unique matrix for each path, we set all such matrices equal to a single learnable parameter $\bm{W}$}, i.e. ${\bm{W}}_{\ell_{1},\ell_{2}}^{\ell}\equiv{\bm{W}}$ for every admissible path $(\ell_{1},\ell_{2},\ell_{3})$. This collapses the $\mathcal{O}(L^{3})$ distinct weight tensors of the naïve implementation into one, reducing the parameter count from $\mathcal{O}\bigl(cL^{3}\bigr)$ to $\mathcal{O}\bigl(c\bigr)$. Because the sharing operates exclusively on multiplicity indices, the irrep content of each block and hence full $\mathrm{SO}(3)$ equivariance is retained. The resulting layer therefore generalizes the classic CG tensor product while offering an order-of-magnitude reduction in parameters. We also extend this scheme to equivariant linear layers and use it in our main experiments.
@@ 3.3 Complexity Analysis of Approximate Tensor Product
In this section, we analyze the computational complexity of the proposed approximate tensor product using CP decomposition. We first investigate the rank of the tensor product. Let $\operatorname{rank}_{\mathrm{CP}}({\bm{M}})$ denote the CP rank of the three-way tensor ${\bm{M}}\in\mathbb{R}^{d_{3}\times d_{1}\times d_{2}}$, representing the minimal rank $R$ such that there is an equality for Eq.˜2, i.e.
@@
By using CP decomposition, the computational cost for evaluating the approximate tensor product in Eq.˜3 reduces to $\mathcal{O}\bigl(R(d_{1}+d_{2}+d_{3})\bigr)$, a significant reduction compared to the full tensor contraction cost of $\mathcal{O}(d_{1}d_{2}d_{3})$ using Eq.˜1. Similarly, the parameter budget decreases from $\mathcal{O}(d_{1}d_{2}d_{3})$ to $\mathcal{O}(R(d_{1}+d_{2}+d_{3}))$, which provides substantial savings when $R$ is small.
For the approximation of the CG tensor product, where $d_{1},d_{2},d_{3}\propto\mathcal{O}(L^{2})$, the computational complexity further reduces to $\mathcal{O}(RL^{2})$. In our experiments, we select $R=7L^{2}$. The error curve for the CP decomposition with varying $R$ is discussed in Section˜4.1.
@@ 3.4 Error Bound and Universality Analysis
The error bound and universality analysis provide theoretical guarantees for the CP decomposition of the tensor product in the proposed Tensor Decomposition Network (TDN). This section establishes the error bound for both the approximation and equivariance, demonstrating how the approximation error depends on the spectral tail of the tensor’s singular values. Additionally, we formalize the universality property of the CP decomposition, showing that as the rank $R$ increases, the CP-decomposition-based tensor product can accurately approximate any $\mathrm{SO}(3)$-equivariant bilinear map, thereby preserving the expressive power of the tensor product while reducing computational complexity.
@@
Given a rank $R\leq\operatorname{rank}_{\mathrm{CP}}({\bm{M}})$, one seeks CP-decomposition-based approximation $\widehat{\bm{M}}$ by minimizing $\lVert{\bm{M}}-\widehat{\bm{M}}\rVert$ in a chosen norm, typically the Frobenius norm. Although the non‑convex optimization may admit spurious local minima, the optimization of CP decomposition possesses an essentially unique best rank‑$R$ approximation under Kruskal’s condition (kolda2009tensor, ), and modern alternating least‑squares or gradient methods converge to it under mild coherence assumptions (yang2023global, ). For error estimation in our setting, we specialize to the CG tensor product, where we let all irreps have the same maximum degree and all dimensions equal such that $d=d_{1}=d_{2}=d_{3}$. Given the singular values $\sigma_{k}^{(n)}$ of the mode-$n$ matricization ${\bm{M}}_{(n)}$ of $\bm{M}$} and truncating each ${\bm{M}}_{(n)}$ to rank $R_{T}=\lceil R^{1/3}\rceil$, a priori approximation error bound (de2000multilinear, ) gives
@@
The following result bounds this error uniformly over $\rm{SO}(3)$ with the proof in Section˜B.1.
Theorem 3.1 (Equivariance Error Bound of CP Decomposition): Let CG tensor ${\bm{M}}\in\mathbb{R}^{d\times d\times d}$ and $\widehat{{\bm{M}}}$ be the rank-$R$ CP-decomposition-based approximation obtained by Frobenius minimization. For any rotation ${\bm{R}}\in SO(3)$ and any bounded irreps ${\bm{x}},{\bm{y}}\in\mathbb{R}^{d},\lVert{\bm{x}}\rVert,\lVert{\bm{y}}\rVert\leq C$, we have $\varepsilon({\bm{R}},{\bm{x}},{\bm{y}})\;\leq\;2C^{2}\Bigl(\sum_{n=1}^{3}\sum_{k>R_{T}}\sigma_{k}^{(n)2}\Bigr)^{1/2},$ where $R_{T}=\lceil R^{1/3}\rceil$ and $\sigma_{k}^{(n)}$ is the $k$-th singular value of mode-$n$ matricization of $\bm{M}$}.
Empirical estimates of both the approximation and equivariance errors are provided in Section˜4.1.
@@
Because the tensor product is universal for bilinear maps, any $\rm{SO}(3)$‑equivariant bilinear operator can be expressed as a composition of a tensor product followed by a suitable $\bm{M}$} onto an equivariant subspace. Consequently, the CP-decomposition‑based approximation developed above inherits this universality: for every equivariant tensor product there exists a rank‑$R$ approximation that converges to it as $R\to\operatorname{rank}_{\mathrm{CP}}({\bm{M}})$. The following theorem formalizes the expressivity of our approximation scheme with the proof in Section˜B.2.
Theorem 3.2 (Universality of CP Decomposition): Consider $\rm{SO}(3)$-irreps ${\bm{x}}\in V_{1}=\mathbb{R}^{d}$, ${\bm{y}}\in V_{2}=\mathbb{R}^{d_{2}}$ and $V_{3}=\mathbb{R}^{d_{3}}$. For any $\rm{SO}(3)$-equivariant bilinear map $\Phi$, there exist ${\bm{B}}\in\mathbb{R}^{d_{1}\times R}$, ${\bm{C}}\in\mathbb{R}^{d_{2}\times R}$, ${\bm{A}}\in\mathbb{R}^{R\times d_{3}}$ such that $\Phi$ can be written as $\Phi({\bm{x}},{\bm{y}})={\bm{A}}({\bm{B}}^{\top}{\bm{x}}\odot{\bm{C}}^{\top}{\bm{y}})\in V_{3},$ if $R$ is sufficiently large.